Shift report 20020120 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 20 2002 - 08:09:59 EST

  • Next message: Stephen J. Sanders: "Re: Shift report 20020119 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Jan 20 08:09:58 2002, Shift: 00:00-08:00
    Supervisor: BHS, with: 
    Log book#12 pages 90 - 93
    
    Shift Summary:
    ------------------------------------------------------------
    Took over a store from last shift, which lasted until a 
    snake quench killed it at 01:20. Another store was achieved
    at 04:35 and is still running (despite desperate 
    protests from PHENIX). Beam will be dumped within the next
    hour and the machine will have a polarimeter calibration run.
    Data runs:
    6691, 6703, 6704, 6706 (running).
    Grand total: 
    50+k triggers in FS at 12deg 
    40k in MRS@90deg, 6B
    32k in MRS@90deg, 2A
    and an added bonus:
    40k in MRS@90deg, 0 field (run 6704)
    because I waited 1hr 15mins for an engineer who kept saying
    he would turn up "within the next five minutes", and when 
    he did come had forgotten his wrench...
    Problems:
    The DAQ lost its connection to the Event builder during the
    night and needed slightly more serious restarting. Thanks to
    JH for the help.
    
    BFS still won't move - I have nudged it now and again to see 
    if it would suddenly wake up. (Well, at least it woke me up...)
    
    Non-problems:
    It's been a while since I've seen such a picture perfect
    winter morning. Two inches of fresh snow, and bright sunshine :-)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Jan 20 2002 - 08:10:33 EST